# 0.5 eV QCD Axion Cyclic Cosmology

### v46

@@ 1 Introduction
Figure 4: Axion-photon coupling grows with initial core helium mass fraction in simulations of stellar evolution for horizontal branch stars constrained by observations of globular clusters (curved line with dots) (A. Ayala et al., 2014). The predicted value for the axion-photon coupling strength (horizontal lines with boxes), given in Eq. (8), combines with this empirical relation to give a prediction for the helium mass fraction (box with whiskers) that differs by $5.0\sigma$ from the primordial helium abundance expected within the standard cold dark matter cosmology with cosmological constant ($\rm\Lambda CDM$}) based on simulations of Big Bang nucleosynthesis (thick vertical line) (S. Gariazzo et al., 2022; C. Pitrou et al., 2018), but that agrees rather well with the best present estimates for the initial core helium mass fraction in the Sun (pair of thin vertical lines) (L. Piersanti et al., 2007; A. M. Serenelli & S. Basu, 2010).
@@
Figure 4 shows the empirical relation between the axion-photon coupling strength, $g_{A\gamma\gamma}=g_{10}\times 10^{-10}~{\rm GeV}^{-1}$, and the initial core helium mass fraction, $Y_{\rm HB}$, for stars that are now on the horizontal branch (HB) of the color-magnitude diagram for globular clusters (A. Ayala et al., 2014):
@@
where $R=N_{\rm HB}/N_{\rm RGB}=1.39~(3)$ was the observed ratio of the number of HB stars, $N_{\rm HB}$, to the number of stars on the red giant branch (RGB) of the color-magnitude diagram, $N_{\rm RGB}$, in 39 globular clusters. Combining the empirical relation in Eq. (34) with the predicted value for the axion-photon coupling in Eq. (8), we find a percent-level precise estimate for the initial core helium mass fraction in these HB stars that formed in the first billion years after the Big Bang:
@@
which differs by $5\sigma$ with the best present estimate of the “primordial” helium abundance, $Y_{P}({\rm\Lambda CDM})=0.24709~(17)~[0.69~{\rm per~mille}]$ (C. Pitrou et al., 2018; S. Gariazzo et al., 2022), within the standard cold dark matter cosmology, but that roughly matches the best present estimates for the initial core helium mass fraction in the Sun, $Y_{\odot}=0.273~(6)~[{\rm 22~per~mille}]$ (A. M. Serenelli & S. Basu, 2010; L. Piersanti et al., 2007). As discussed in Section 6, the mix of axions and photons that dominates the expansion of the universe leading up to Big Bang nucleosynthesis, within the proposed physical picture, yields about ten percent more primordial helium, $Y_{P}\approx 0.27$, than the standard mix of light leptons and photons, and, in this way, resolves the $5\sigma$ tension between $Y_{\rm HB}$ in Eq. (35) and $Y_{P}({\rm\Lambda CDM})$.
Similarly, the recent attempts to “rule out” axion dark matter with the rest-mass energy and axion-photon coupling indicated by laboratory astrophysics (See Section 8), using “blank-sky” observations of the dark matter “halo” of the Milky Way at the near-infrared wavelength $2\lambda_{A}=2h/m_{A}c\approx 4.93~{\rm\mu m}$, corresponding to the decay of the axion into a pair of photons, subtract a “smooth astrophysical background” by fitting a quartic polynomial to the near-infrared spectrum within a window of width $\Delta\lambda\approx 0.2~{\rm\mu m}$ (E. Pinetti, 2025). However, the proposed physical picture implies that 0.5 eV QCD axion dark matter is in thermal equilibrium with the cosmic microwave background (See Section 3 and Appendix A), and, thus, the axions in the dark matter “halo” of the Milky Way have velocity dispersion, $\sigma_{v}$, fixed by the CMB temperature, $T_{\gamma}$, and by the 0.5 eV QCD axion dark matter rest-mass energy, $m_{A}c^{2}$, leading to an expected line-width, $2\Delta\lambda_{A}$, that is roughly the same size as the window used to subtract the “smooth astrophysical background:”
@@
Turning to axion-electron coupling, the key point is that there is no “tree-level” or “bare” coupling between the axion and electron R. D. Peccei (1989), just as there is no such coupling between the neutral pion, $\pi^{0}$, and the electron (T. Husek, 2024). Instead, the “leading order” axion-electron coupling appears only at “loop-level,” and is suppressed by the small factor, $(\frac{\alpha}{2\pi})^{2}\sim 10^{-6}$ (P. J. Mohr et al., 2025), compared to the “bare” value used to derive astrophysical “constraints” and “exclusions” (K. Sato & H. Sato, 1975; K. Sato, 1978; A. Caputo & G. Raffelt, 2024). The upshot is that the effect of axions on the astrophysical phenomena, such as white dwarf cooling and red giant luminosity, is smaller than the astrophysical uncertainties entering the analysis, and, thus, no significant “constraints” or “exclusions” emerge from the comparison with astrophysical observations.
Finally, regarding the axion-nucleon coupling, a dense bunch of nucleons makes a lot of 0.5 eV QCD axions due to this coupling, but, by the same token, such a dense nuclear soup is “axionically thick,” in the sense that a 0.5 eV QCD axion made by one nucleon is unlikely to “escape” the concentration without coupling to another nearby nucleon, as pointed out by M. S. Turner (1988), in the analysis of the effect of axions on supernova 1987A. Remarkably, there is a relatively narrow window of QCD axion rest-mass energy, $0.05~{\rm eV}\leq m_{A}c^{2}\leq 5~{\rm eV}$ (W. Yin, 2023, pg. 42), that brackets the predicted value in Eq. (7), in which the effects of the QCD axions completely dominate the flow of energy within the dense nuclear soup, in sharp contrast with the traditional assumption that axions play essentially no role in this flow, and this contrast is often expressed as an “exclusion” of the 0.5 eV QCD axion. However, given the direct laboratory astrophysics evidence for 0.5 eV QCD axion dark matter, sketched in Section 8, one is instead led, on the basis of this analysis of the axion-nucleon coupling, to question “astrophysicists’ most cherished beliefs” (A. Burrows et al., 1989) regarding the flow of energy within compact astrophysical bodies, such as pulsars, and within the “core-collapse” supernovae that are thought to form such bodies.
@@ 10 Conclusions
Moving forward, the sharp difference in the baryon abundance, roughly $180\sigma$, gives about a factor of two more baryons with the present physical picture. At the same time, the mix of six axions for every photon made in the Big Bang, within this picture, speeds up the cooling off of the universe in the first three minutes. The combination of these two sharp differences leads to about the same deuterium abundance but about $5\sigma$ more helium compared to $\rm\Lambda CDM$}, we have argued, and these loose arguments, in part based on astrophysical observations of globular clusters, must be confirmed by actual BBN calculations within the proposed physical picture.
In a similar vein, an actual calculation of the spectrum of CMB fluctuations within the proposed physical picture must be done. We have argued, indirectly, that the location of the main peak in the spectrum should come out about the same as it does in $\rm\Lambda CDM$} due to the coincidence of the redshift where the 0.5 eV QCD axion dark matter equation of state crosses over from radiation-like to matter-like, on the one hand, and, on the other, the redshift of “matter-radiation equality” within $\rm\Lambda CDM$}. Along similar lines, there is reason to believe that the sharp difference in baryon abundance, roughly $180\sigma$, combines with the sharp difference in the “primordial” spectral index, nearly $9\sigma$, to give basically the same spectrum as the best fit within $\rm\Lambda CDM$}, so that the agreement of this best fit with early universe observations is retained in the present physical picture, albeit with first-principles predictions taking the place of the key fit parameters.
Finally, and perhaps most importantly, the direct laboratory astrophysics evidence for the creation of 0.5 eV QCD axion dark matter in a straightforward, “table-top” experiment, presented in this Letter, suggests that we must “re-visit” many of the conventional assumptions about the way dark matter interacts with light energy and “bright” matter. For much of the past half century, it has been conventional to assume that dark matter “tracks” bright matter in gravitational equilibrium within a “halo” around the Milky Way galaxy, but the direct laboratory astrophysics evidence for 0.5 eV QCD axion dark matter, when combined with the lack of a sharp two-photon decay “line” from 0.5 eV QCD axion dark matter in the infrared spectrum of this halo, suggests that dark matter is, instead, in thermal equilibrium with the photons of the CMB, as predicted by the proposed physical picture. Even more iconoclastically, the direct laboratory evidence for 0.5 eV QCD axion dark matter, when combined with long-standing analyses of the effects of axion-nucleon coupling on dense nuclear soup in astrophysical settings, suggests that “astrophysicists’ most cherished beliefs” about core-collapse supernovae and the internal structure of pulsars are simply wrong, and that, instead, 0.5 eV QCD axion dark matter completely dominates the flow of energy within compact astrophysical bodies and within the events that form them.
@@
At the Big Bang, the average kinetic energy, $\check{E}=x_{p}kT_{\pi}\approx 104~{\rm MeV}$, given in Eq. (17), is much larger than the 0.5 eV QCD axion dark matter rest-mass energy, $m_{A}c^{2}=0.5~{\rm eV}=5\times 10^{-7}~{\rm MeV}$, given in Eq. (7). The axion number density at the Big Bang, $n_{A}(T_{\pi})$, is then determined by the Planck distribution function, just as the photon number density at the Big Bang, $n_{\gamma}(T_{\pi})$, is so determined (P. J. E. Peebles, 1993, pg. 134–137, 158–159), with the result
@@
where $n_{A}$ is the axion number density in the present universe, $T_{\gamma}$ is the temperature of the cosmic microwave background, and the statistical degeneracy factor, $g_{A}=12$, counts the twelve “flavors” of quarks and leptons in the standard model of particle physics (L. M. Lederman & D. N. Schramm, 1989, pg. 128). Each flavor, $M$, of quark and lepton binds with an anti-particle of the same flavor and helicity to make the axion, $A_{M}$, with this flavor and no spin, such that the axion number density at the Big Bang is given by the expression in Eq. (10), with the value of the dimensionless constant, $N_{n}=6\times 2!\zeta(3)/(\pi^{2}x_{p}^{3})$, fixed by the result given above in Eq. (A3).
@@
The predicted value for the rest-mass energy, $m_{A}c^{2}$, of the 0.5 eV QCD axion dark matter, given in Eq. (7), follows from the analysis of the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, presented at the end of Section 3, once a relation can be found with the temperature of the Big Bang, $T_{\pi}$. Both of the quantities, $m_{A}c^{2}$ and $kT_{\pi}$, enter the statistical mechanical expression for this ratio, given in Eq. (25), through their ratios with the nucleon rest-mass energies, $m_{N}c^{2}$, for N=n,pN=n,~p. Given an independent relation between $m_{A}c^{2}$ and $kT_{\pi}$, then, the matching of this expression to the numerically exact value for the ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}=2(x_{p}y_{T})^{2}/\pi^{2}$, given in Eq. (5) and derived in Section 4, determines both quantities with a precision that is limited only by that of this independent relation.
The starting point for finding such an independent relation between $m_{A}c^{2}$ and $kT_{\pi}$ is the relation between the axion decay constant, $f_{A}$, and the kinetic energy, $\hat{E}$, of the massless particle and anti-particle within the axion at future infinity, given in Eq. (14), combined with the “see-saw” relation for the value of this kinetic energy, $\check{E}=x_{p}kT_{\pi}$, at the Big Bang, given in Eq. (13):
$f_{A}$ $=$ $\left(\frac{f_{A}^{2}}{x_{p}kT_{\pi}}\right)^{3}\left(E_{\rm Pl}\right)^{-2}$ (B1)
$=$ $f_{A}\left(\frac{\sqrt{\chi_{\rm QCD}}}{m_{A}c^{2}}\right)^{5}\left(x_{p}kT_{\pi}\right)^{-3}\left(E_{\rm Pl}\right)^{-2}$
$=$ $f_{A}\frac{1}{x_{p}^{3}y_{T}^{10}}\left(kT_{\pi}\right)^{7}\left(m_{A}c^{2}\right)^{-5}\left(E_{\rm Pl}\right)^{-2}$
@@
$m_{A}c^{2}=x_{p}^{-3/5}y_{T}^{-2}\left(kT_{\pi}\right)^{7/5}\left(E_{\rm Pl}\right)^{-2/5}.$ (B2)
The precision of this relation is set by the 22 ppm uncertainty in the best present value for the universal gravitational constant, $G=6.674~08~(15)\times 10^{-8}~{\rm g^{-1}~cm^{3}~s^{-2}}$ (P. J. Mohr et al., 2025), and we find the relative uncertainties, $\Delta X/X=|n_{X}|\Delta G/G$, for the predicted values of various quantities, $X$, in terms of the relative uncertainty of $G$ and the power, $n_{X}$, of the dependence, $X\propto G^{n_{X}}$:
$\frac{\Delta m_{A}c^{2}}{m_{A}c^{2}}=\frac{1}{5}\frac{\Delta G}{G}={\rm 4.4~ppm}$
$\frac{\Delta kT_{\pi}}{kT_{\pi}}=\frac{1}{7}\frac{\Delta G}{G}={\rm 3.2~ppm}$
$\frac{\Delta f_{A}}{f_{A}}=\frac{3}{35}\frac{\Delta G}{G}={\rm 1.9~ppm}$
$\frac{\Delta\chi_{\rm QCD}}{\chi_{\rm QCD}}=\frac{4}{7}\frac{\Delta G}{G}={\rm 13~ppm}$
$\frac{\Delta\Omega_{\Lambda}h^{2}}{\Omega_{\Lambda}h^{2}}=2\frac{\Delta G}{G}={\rm 44~ppm}.$ (B3)
@@
To make the connection, we use the field equations of general relativity to express the value of the Ricci curvature scalar at future infinity, $\hat{R}$, in terms of the cosmological constant, $\Lambda$, and the value at the Big Bang, $\check{R}$, in terms of the baryon energy density at the Big Bang (R. Penrose & W. Rindler, 1986, pg. 20)
@@
$\check{R}$ $=$ $\frac{8\pi G}{c^{4}}\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}u_{B},$ (C3)
@@
$4\Omega_{\Lambda}h^{2}\left(\frac{\hat{E}}{f_{A}}\right)^{2}=\Omega_{B}h^{2}\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}\left(\frac{\check{E}}{f_{A}}\right)^{2},$ (C4)
@@
$=$ $1.~047~007~(2)\times 10^{-13}~{\rm MeV}^{-1}~C_{A\gamma}~[1.9~{\rm ppm}],$
where $\alpha=e^{2}/(\hbar c)=1/(137.~035~999~18~(2))$ (P. J. Mohr et al., 2025) [1.5 parts per billion] is the best present estimate for the dimensionless electromagnetic “fine-structure” coupling constant, and where we have used, in the second line, the relation for the QCD axion decay constant, $f_{A}=\sqrt{\chi_{\rm QCD}}/(m_{A}c^{2})$ (S. Weinberg, 1978; F. Wilczek, 1978), together with the values for the axion rest-mass energy, $m_{A}c^{2}$, given in Eq. (7), and the QCD topological susceptibility, $\chi_{\rm QCD}$, given by Eq. (3) (See Appendix B for the estimate of the uncertainty in the value of $f_{A}$). To estimate the value of $C_{A\gamma}$, we use its relation with the light-quark mass ratio, $m_{u}/m_{d}=0.485~(19)$ (Z. Fodor et al., 2016), and with the electromagnetic-to-strong axial anomaly ratio, $E/N=8/3$, (R. D. Peccei, 1989; M. Gorghetto & G. Villadoro, 2019),

### v47

@@ 1 Introduction
Figure 3: Axion-photon coupling grows with initial core helium mass fraction in simulations of stellar evolution for horizontal branch stars constrained by observations of globular clusters (curved line with dots) (A. Ayala et al., 2014). The predicted value for the axion-photon coupling strength (horizontal lines with boxes), given in Eq. (7), combines with this empirical-computational relation to give a prediction for the helium mass fraction (box with whiskers) that differs by $5.0\sigma$ from the primordial helium abundance expected within the standard cold dark matter cosmology with cosmological constant ($\rm\Lambda CDM$}) based on simulations of Big Bang nucleosynthesis (thick vertical line) (S. Gariazzo et al., 2022; C. Pitrou et al., 2018), but that agrees rather well with the best present estimates for the initial core helium mass fraction in the Sun (pair of thin vertical lines) (L. Piersanti et al., 2007; A. M. Serenelli & S. Basu, 2010).
@@
Figure 4 shows the empirical-computational relation between the axion-photon coupling strength, $g_{A\gamma\gamma}=g_{10}\times 10^{-10}~{\rm GeV}^{-1}$, and the initial core helium mass fraction, $Y_{\rm HB}$, for stars that are now on the horizontal branch (HB) of the color-magnitude diagram for globular clusters (A. Ayala et al., 2014):
@@
where $R=N_{\rm HB}/N_{\rm RGB}=1.39~(3)$ was the observed ratio of the number of HB stars, $N_{\rm HB}$, to the number of stars on the red giant branch (RGB) of the color-magnitude diagram, $N_{\rm RGB}$, in 39 globular clusters. Combining the empirical-computational relation in Eq. (32) with the predicted value for the axion-photon coupling in Eq. (7), we find a percent-level precise estimate for the initial core helium mass fraction in these HB stars that formed in the first billion years after the Big Bang:
@@
which differs by $5\sigma$ with the best present estimate of the “primordial” helium abundance, $Y_{P}({\rm\Lambda CDM})=0.247~09~(17)~[0.69~{\rm per~mille}]$ (C. Pitrou et al., 2018; S. Gariazzo et al., 2022), within the standard cold dark matter cosmology, but that roughly matches the best present estimates for the initial core helium mass fraction in the Sun, $Y_{\odot}=0.273~(6)~[{\rm 22~per~mille}]$ (A. M. Serenelli & S. Basu, 2010; L. Piersanti et al., 2007). As discussed in Section 6, the mix of axions and photons that dominates the expansion of the universe leading up to Big Bang nucleosynthesis, within the proposed physical picture, yields about ten percent more primordial helium, $Y_{P}\approx 0.27$, than the standard mix of light leptons and photons, and, in this way, resolves the $5\sigma$ tension between $Y_{\rm HB}$ in Eq. (33) and $Y_{P}({\rm\Lambda CDM})$.
Similarly, the recent attempts to “rule out” axion dark matter with the rest-mass energy in Eq. (6) and the axion-photon coupling in Eq. (7) using “blank-sky” observations of the dark matter “halo” of the Milky Way at the near-infrared wavelength $2\lambda_{A}=2h/m_{A}c\approx 4.92~{\rm\mu m}$, corresponding to the decay of the axion into a pair of photons, subtract a “smooth astrophysical background” by fitting a quartic polynomial to the near-infrared spectrum within a window of width $\Delta\lambda\approx 0.2~{\rm\mu m}$ (E. Pinetti, 2025). However, the proposed physical picture implies that 0.5 eV QCD axion dark matter is in thermal equilibrium with the cosmic microwave background (See Section 3 and Appendix A), and, thus, the axions in the dark matter “halo” of the Milky Way have velocity dispersion, $\sigma_{v}$, fixed by the CMB temperature, $T_{\gamma}$, and by the 0.5 eV QCD axion dark matter rest-mass energy, $m_{A}c^{2}$, leading to an expected line-width, $2\Delta\lambda_{A}$, that is roughly the same size as the window used to subtract the “smooth astrophysical background:”
@@
Turning to axion-electron coupling, the key point is that there is no “tree-level” or “bare” coupling between the axion and electron (R. D. Peccei, 1989), just as there is no such coupling between the neutral pion, $\pi^{0}$, and the electron (T. Husek, 2024). Instead, the “leading order” axion-electron coupling appears only at “loop-level,” and is suppressed by the small factor, $(\frac{\alpha}{2\pi})^{2}\sim 10^{-6}$ (P. J. Mohr et al., 2025), compared to the “bare” value used to derive astrophysical “constraints” and “exclusions” (K. Sato & H. Sato, 1975; K. Sato, 1978; A. Caputo & G. Raffelt, 2024). The upshot is that the effect of axions on the astrophysical phenomena, such as white dwarf cooling and red giant luminosity, is smaller than the astrophysical uncertainties entering the analysis, and, thus, no significant “constraints” or “exclusions” emerge from the comparison with astrophysical observations.
Finally, regarding the axion-nucleon coupling, a dense bunch of nucleons makes a lot of 0.5 eV QCD axions due to this coupling, but, by the same token, such a dense nuclear soup is “axionically thick,” in the sense that a 0.5 eV QCD axion made by one nucleon is unlikely to “escape” the concentration without coupling to another nearby nucleon, as pointed out by M. S. Turner (1988), in the analysis of the effect of axions on supernova 1987A. Remarkably, there is a relatively narrow window of QCD axion rest-mass energy, $0.05~{\rm eV}\leq m_{A}c^{2}\leq 5~{\rm eV}$ (W. Yin, 2023, pg. 42), that brackets the predicted value in Eq. (6), in which the effects of the QCD axions completely dominate the flow of energy within the dense nuclear soup, in sharp contrast with the traditional assumption that axions play essentially no role in this flow, and this contrast is often expressed as an “exclusion” of the 0.5 eV QCD axion. However, given the apparent success of 0.5 eV QCD axion dark matter in resolving the Hubble tension, which is the main focus of this Letter, one is instead led, on the basis of this analysis of the axion-nucleon coupling, to question the traditional assumptions regarding the flow of energy within compact astrophysical bodies, such as pulsars, as well as the flow within the events that are thought to form them, such as Type II supernovae.
@@ 9 Conclusions
Moving forward, the sharp difference in the baryon abundance, roughly $190\sigma$, gives about a factor of two more baryons with the present physical picture. At the same time, the mix of six axions for every photon made in the Big Bang, within this picture, speeds up the cooling off of the universe in the first three minutes. The combination of these two sharp differences leads to about the same deuterium abundance but about $5\sigma$ more helium compared to $\rm\Lambda CDM$}, we have argued, and these loose arguments, in part based on astrophysical observations of globular clusters, must be confirmed by actual BBN calculations within the proposed physical picture.
In a similar vein, an actual calculation of the spectrum of CMB fluctuations within the proposed physical picture must be done. We have argued, indirectly, that the location of the main peak in the spectrum should come out about the same as it does in $\rm\Lambda CDM$} due to the coincidence of the redshift where the 0.5 eV QCD axion dark matter equation of state crosses over from radiation-like to matter-like, on the one hand, and, on the other, the redshift of “matter-radiation equality” within $\rm\Lambda CDM$}. Along similar lines, there is reason to believe that the sharp difference in baryon abundance, roughly $190\sigma$, combines with the sharp difference in the “primordial” spectral index, nearly $9\sigma$, to give basically the same spectrum as the best fit within $\rm\Lambda CDM$}, so that the agreement of this best fit with early universe observations is retained in the present physical picture, albeit with first-principles predictions taking the place of the key fit parameters.
Finally, the straightforward resolution of the Hubble tension using 0.5 eV QCD axion cyclic cosmology, presented in this Letter, suggests that we must “re-visit” many of the conventional assumptions about the way dark matter interacts with light energy and “bright” matter. For much of the past half century, it has been conventional to assume that dark matter “tracks” bright matter in gravitational equilibrium within a “halo” around the Milky Way galaxy, but the lack of a sharp two-photon decay “line” from 0.5 eV QCD axion dark matter in the infrared spectrum of this halo, suggests that dark matter is, instead, in thermal equilibrium with the photons of the CMB, as predicted by the proposed physical picture. For an equally long span of time, axions have been assumed to play essentially no role in the flow of energy within compact astrophysical bodies, such as pulsars, and within Type II supernovae, such as supernova 1987A, and, thus, it may be time to re-visit those assumptions, in light of the resolution of the Hubble tension presented in this Letter.
@@
At the Big Bang, the average kinetic energy, $\check{E}=x_{k}kT_{\pi}\approx 104~{\rm MeV}$, given in Eq. (16), is much larger than the 0.5 eV QCD axion dark matter rest-mass energy, $m_{A}c^{2}=0.5~{\rm eV}=5\times 10^{-7}~{\rm MeV}$, given in Eq. (6). The axion number density at the Big Bang, $n_{A}(T_{\pi})$, is then determined by the Planck distribution function, just as the photon number density at the Big Bang, $n_{\gamma}(T_{\pi})$, is so determined (P. J. E. Peebles, 1993, pg. 134–137, 158–159), with the result
@@
where $n_{A}$ is the axion number density in the present universe, $T_{\gamma}$ is the temperature of the cosmic microwave background, and the statistical degeneracy factor, $g_{A}=12$, counts the twelve “flavors” of quarks and leptons in the standard model of particle physics (L. M. Lederman & D. N. Schramm, 1989, pg. 128). Each flavor, $M$, of quark and lepton binds with an anti-particle of the same flavor and helicity to make the axion, $A_{M}$, with this flavor and no spin, such that the axion number density at the Big Bang is given by the expression in Eq. (9), with the value of the dimensionless constant, $N_{n}=6\times 2!\zeta(3)/(\pi^{2}x_{k}^{3})$, fixed by the result given above in Eq. (A3).
@@
The predicted value for the rest-mass energy, $m_{A}c^{2}$, of the 0.5 eV QCD axion dark matter, given in Eq. (6), follows from the analysis of the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, presented at the end of Section 3, once a relation can be found with the temperature of the Big Bang, $T_{\pi}$. Both of the quantities, $m_{A}c^{2}$ and $kT_{\pi}$, enter the statistical mechanical expression for this ratio, given in Eq. (24), through their ratios with the nucleon rest-mass energies, $m_{N}c^{2}$, for N=n,pN=n,~p. Given an independent relation between $m_{A}c^{2}$ and $kT_{\pi}$, then, the matching of this expression to the numerically exact value for the ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, given in Eq. (5) and derived in Section 4, determines both quantities with a precision that is limited only by that of this independent relation.
The starting point for finding such an independent relation between $m_{A}c^{2}$ and $kT_{\pi}$ is the relation between the axion decay constant, $f_{A}$, and the kinetic energy, $\hat{E}$, of the massless particle and anti-particle within the axion at future infinity, given in Eq. (13), combined with the “see-saw” relation for the value of this kinetic energy, $\check{E}=x_{k}kT_{\pi}$, at the Big Bang, given in Eq. (12):
$f_{A}$ $=$ $\left(\frac{f_{A}^{2}}{x_{k}kT_{\pi}}\right)^{3}\left(E_{\rm Pl}\right)^{-2}$ (B1)
$=$ $f_{A}\left(\frac{\sqrt{\chi_{\rm QCD}}}{m_{A}c^{2}}\right)^{5}\left(x_{k}kT_{\pi}\right)^{-3}\left(E_{\rm Pl}\right)^{-2}$
$=$ $f_{A}\frac{1}{x_{k}^{3}y_{T}^{10}}\left(kT_{\pi}\right)^{7}\left(m_{A}c^{2}\right)^{-5}\left(E_{\rm Pl}\right)^{-2}$
@@
$m_{A}c^{2}=x_{k}^{-3/5}y_{T}^{-2}\left(kT_{\pi}\right)^{7/5}\left(E_{\rm Pl}\right)^{-2/5}.$ (B2)
The precision of this relation is set by the 22 ppm uncertainty in the best present value for the universal gravitational constant, $G=6.674~08~(15)\times 10^{-8}~{\rm g^{-1}~cm^{3}~s^{-2}}$ (P. J. Mohr et al., 2025), and, we obtain the precision of the determinations of $m_{A}c^{2}$ and $kT_{\pi}$ by simply varying the value of $E_{\rm Pl}$ in Eq. (B2), within the 11 ppm window set by $\Delta G/G$, and reading off the changes, $\Delta m_{A}c^{2}$ and $\Delta kT_{\pi}$, with the following empirical-computational results:
$\frac{\Delta m_{A}c^{2}}{m_{A}c^{2}}$ $=$ $24\times 10^{-3}~\frac{\Delta G}{G}={\rm 0.53~ppm}$
$\frac{\Delta kT_{\pi}}{kT_{\pi}}$ $=$ $7.7\times 10^{-3}~\frac{\Delta G}{G}={\rm 0.17~ppm}$
$\frac{\Delta\chi_{\rm QCD}}{\chi_{\rm QCD}}$ $=$ $4\frac{\Delta kT_{\pi}}{kT_{\pi}}={\rm 0.68~ppm}$
$\frac{\Delta f_{A}}{f_{A}}$ $=$ $\left|2\frac{\Delta kT_{\pi}}{kT_{\pi}}-\frac{\Delta m_{A}c^{2}}{m_{A}c^{2}}\right|={\rm 0.19~ppm}$
$\frac{\Delta\Omega_{\Lambda}h^{2}}{\Omega_{\Lambda}h^{2}}$ $=$ $\frac{\Delta G}{G}={\rm 22~ppm},$ (B3)
where the precision of the derived quantities, $\chi_{\rm QCD}=y_{T}^{-4}(kT_{\pi})^{4}$, $f_{A}=\sqrt{\chi_{\rm QCD}}/(m_{A}c^{2})=y_{T}^{-2}(kT_{\pi})^{2}(m_{A}c^{2})^{-1}$, and, $\Omega_{\Lambda}h^{2}\propto G$, follows from straight-forward error-propagation (P. J. Mohr et al., 2025).
@@
To make the connection, we use the field equations of general relativity to express the value of the Ricci curvature scalar at future infinity, $\hat{R}$, in terms of the cosmological constant, $\Lambda$, and the value at the Big Bang, $\check{R}$, in terms of the trace of the baryon stress-energy tensor at the Big Bang (R. Penrose & W. Rindler, 1986, pg. 20)
@@
$\check{R}$ $=$ $\frac{8\pi G}{c^{4}}\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}\left(\frac{1+\epsilon}{1+\delta}\right)u_{B},$ (C3)
@@
$4\Omega_{\Lambda}h^{2}\left(\frac{\hat{E}}{f_{A}}\right)^{2}=\Omega_{B}h^{2}\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}\left(\frac{\check{E}}{f_{A}}\right)^{2}\left(\frac{1+\epsilon}{1+\delta}\right),$ (C4)
@@
We estimate the small dimensionless quantity, $\epsilon=(4m_{\alpha}/m_{p}-1)Y_{P}=1.876~(34)\times 10^{-3}~[18~{\rm per~mille}]$, using the best present value for the ratio, $m_{\alpha}/m_{p}=3.972~599~690~252~(70)~[18~{\rm parts~per~trillion}]$ (P. J. Mohr et al., 2025), of the mass of the helium nucleus to the proton mass, and using the astrophysical determination of the primordial helium abundance within the present physical picture, $Y_{P}=Y_{\rm HB}=0.272~(5)~[18~{\rm per~mille}]$ (See Section 8). The exact expression for the relativistic correction, $1+\delta=m_{p}\eta/(m_{p}e^{-x_{p}}x_{p}J(x_{p})+m_{n}e^{-x_{n}}x_{n}J(x_{n}))$, uses the nucleon mass-to-temperature ratios at the Big Bang, $x_{N}=m_{N}c^{2}/kT_{\pi}$, for $N=p$ and $n$, with $kT_{\pi}$ from Eq. (3). The numerically exact integral, $J(x_{N})$, is defined in Eq. (25), while the integral, $I(x_{N})=(1/2!\zeta(3))\int_{0}^{\infty}dxe^{-x}(x+x_{N})\sqrt{x(x+2x_{N})}$, determines the baryon-to-photon number density ratio, $\eta=e^{-x_{p}}I(x_{p})+e^{-x_{n}}I(x_{n})=1.290~113~(5)\times 10^{-9}~[{\rm 4~ppm}].$
@@
$=$ $1.048~854~09~(20)\times 10^{-13}~{\rm MeV}^{-1}~C_{A\gamma}~[0.19~{\rm ppm}],$
where $\alpha=e^{2}/(\hbar c)=1/(137.035~999~18~(2))$ (P. J. Mohr et al., 2025) [1.5 parts per billion] is the best present estimate for the dimensionless electromagnetic “fine-structure” coupling constant, and where we have used, in the second line, the relation for the QCD axion decay constant, $f_{A}=\sqrt{\chi_{\rm QCD}}/(m_{A}c^{2})$ (S. Weinberg, 1978; F. Wilczek, 1978), together with the values for the axion rest-mass energy, $m_{A}c^{2}$, given in Eq. (6), and the QCD topological susceptibility, $\chi_{\rm QCD}$, given by Eq. (3) (See Appendix B for the estimate of the uncertainty in the value of $f_{A}$). To estimate the value of $C_{A\gamma}$, we use its relation with the light-quark mass ratio, $m_{u}/m_{d}=0.485~(19)$ (Z. Fodor et al., 2016), and with the electromagnetic-to-strong axial anomaly ratio, $E/N=8/3$, (R. D. Peccei, 1989; M. Gorghetto & G. Villadoro, 2019),